8) Find the average rate of change for the function f(x) = -x2 + 6x on the interval [3,5).
Answer #1

f(b) - f(a) Average = %3D a=3, b > 5 f(5)= -5+6(5) f(3) = - 3*+6(3) = %3D 5-9 Average 5-3 :4